# Deep DeePC: Data-enabled predictive control with low or no online optimization using deep learning

### v2

@@ 1 Introduction
The Koopman theory-based framework [16] holds great promise for data-driven modeling and control of nonlinear systems. By constructing a linear model within a lifted state-space for a general nonlinear system from offline data, Koopman modeling facilitates the application of linear control theory to nonlinear systems and processes. Notable advancements include Koopman-based model predictive control (MPC) approaches [17, 18, 20, 19, 21], which develop linear MPC schemes for nonlinear systems, thereby maintaining convex online optimization despite the inherent nonlinearity of the underlying systems and processes. Koopman-based modeling and control have been widely applied across various fields, including industrial processes [22, 23, 24, 25], vehicles and robotics [26, 27, 28], and power systems [29, 30, 31]. The development of Koopman MPC typically requires full-state feedback from the systems, with exceptions in [32, 33]. In [32], an input-output Koopman control method was proposed, where the Koopman-based controller only requires measurements of partial state variables directly used for calculating operational costs. In [33], a non-exact Koopman multi-step Koopman predictor was identified using only input and output data for the predictive control of nonlinear systems. We note that, in practical applications, measuring certain state variables can be challenging or costly. The need for full-state measurements poses significant challenges to the broader application of Koopman-based control methods in complex industrial systems.
@@
In recent years, data-enabled predictive control (DeePC) has received increasing attention for its ability to address constrained optimal control using only input and output data [36, 34, 35]. DeePC leverages Willems’ fundamental lemma [37] to form a non-parametric representation of the system using a Hankel matrix established based on pre-collected input/output trajectories. Willems’ fundamental lemma indicates that any input/output trajectories of a linear system are spanned by the Hankel matrix when the pre-collected input trajectory is persistently exciting [37]. DeePC produces the optimal control action by solving an optimization problem in a receding-horizon manner, which is similar to MPC [38].
DeePC utilizes gathered data to construct the Hankel matrix, with the dimension of the decision variable in the DeePC optimization problem being related to the dimension of this matrix. To satisfy the persistently exciting condition, a large amount of data is typically required to establish an accurate non-parametric representation of the system. Consequently, as the amount of used data increases, the computational complexity of the DeePC problem also increases, posing challenges for real-time application.
@@
In the existing literature, efforts have been focused on addressing the computational challenges of DeePC by reducing the size or complexity of the Hankel matrix and optimization problems. In [39], singular value decomposition was leveraged on collected data to reduce the dimensionality of the DeePC problem, thus decreasing computational complexity. In [40], proper orthogonal decomposition was implemented to formulate a reduced order control problem, which was subsequently solved using a customized MPC solver. In [41, 42], LQ factorization was employed to redefine the decision variable, replacing the Hankel matrix with a lower triangular matrix. In [43], neighboring extremal was utilized to approximate the DeePC policy and reduce the computational cost for the constrained nonlinear systems. In [44], an online DeePC method was introduced using the primal-dual algorithm to solve the control problem recursively, coupled with the fast Fourier transform to expedite the computation of Hankel matrix-vector products. In [45], an alternative approach was proposed where DeePC is approximated by decomposing the original problem into two parts: a control objective and a scoring function. The scoring function is learned recursively from a size-invariant differentiable convex problem. While these existing methods improve the computational feasibility of DeePC, they still involve solving optimization problems or iterating algorithms during real-time applications, highlighting ongoing challenges in achieving real-time efficiency.
@@
    We propose a deep learning-enabled DeePC method, referred to as Deep DeePC; this method does not need to solve any online optimization when constraints are either absent or not violated.
@@
    We use two benchmark nonlinear process examples to evaluate the performance of the proposed approach. Good control performance is achieved, and the computation time is significantly reduced as compared to conventional DeePC.
2 Preliminaries and problem formulation
2.1 Notation
$\mathbb{R}$ denotes the set of real numbers. $\mathbb{Z}_{>0}$ and $\mathbb{Z}_{\geq 0}$ denote the set of positive integers and the set of non-negative integers, respectively. $\mathbb{E}$ denotes the expectation. $\|x\|_{Q}^{2}$ is the square of the weighted Euclidean norm of vector $x$ with positive-definite weighting matrix $Q$, computed as $\|x\|_{Q}^{2}:=x^{\top}Qx$. $\text{diag}(\cdot)$ denotes a diagonal matrix. $I_{n}$ is an identity matrix of dimension $n$. $\odot$ denotes the Hadamard product. $\mathbf{1}_{a<b}$ denotes an indicator function; it equals $1$ if $a<b$, and equals 00 otherwise. $\mathcal{U}(a,b)$ denotes a uniform distribution with lower and upper bounds $a$ and $b$, respectively. $x(i)$ denotes the $i$th variable of state vector $x$. $x_{k}\in\mathbb{R}^{n_{x}}$ is the state vector at time instant $k$, and $\{x\}_{j}^{l}:=[x_{j}^{\top},\ldots,x_{l}^{\top}]^{\top}$ contains the state sequence from time instant $j$ to $l$.
2.2 Non-parametric representation of linear systems
@@
Let $L$, $T\in\mathbb{Z}_{>0}$ and $T\geq L$. Let $\mathbf{u}_{T}^{d}:=\{u^{d}\}_{1}^{T}\in\mathbb{R}^{n_{u}T}$ and $\mathbf{y}_{T}^{d}:=\{y^{d}\}_{1}^{T}\in\mathbb{R}^{n_{y}T}$ denote the input and output sequences for $T$ time instants, respectively, with the superscript $d$ indicating that the sequences are collected offline. For the input sequence $\mathbf{u}_{T}^{d}$, the Hankel matrix of depth $L$ is defined as follows:
@@
2.3 Data-enabled predictive control (DeePC)
@@ 2.4 Problem formulation
$x_{k+1}$ $=f(x_{k},u_{k})$ (7a)
$y_{k}$ $=Cx_{k}$ (7b)
where $x_{k}\in\mathbb{X}\subset\mathbb{R}^{n_{x}}$ is the system state vector; $u_{k}\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$ is the control input vector; $y_{k}\in\mathbb{Y}\subset\mathbb{R}^{n_{y}}$ is the output measurement vector; $f:\mathbb{X}\times\mathbb{U}\to\mathbb{X}$ is a nonlinear function that characterizes the state dynamics of the process; $C\in\mathbb{R}^{n_{y}\times n_{x}}$ is the system matrix; and $\mathbb{X}$, $\mathbb{U}$, and $\mathbb{Y}$ are compact sets. We consider nonlinear systems in which the output measurements linearly depend on the state vectors. Our objective is to address the control for the nonlinear processes using only input and output data in an efficient manner within the DeePC framework.
@@
3 Applicability of DeePC for nonlinear systems
We will use the conventional DeePC described in (6) as the foundation of developing our deep learning-enabled DeePC method. Given that (6) was originally developed for LTI systems based on Lemma 1, in this section, we leverage the concept of Koopman modeling for control to justify the appropriateness of using conventional DeePC in (6) as the basis for developing DeePC-based controllers for nonlinear systems in (7).
Consider a discrete-time nonlinear control system described in (7). Based on the Koopman theory for controlled systems, the state vector can be extended to include control inputs, that is, $\chi_{k}=[x_{k}^{\top},u_{k}^{\top}]^{\top}$ [17]. Accordingly, there exists an infinite-dimensional nonlinear lifting mapping $\Psi_{\chi}$ that allows the dynamics of the lifted state to be governed by a Koopman operator $\mathcal{K}$ as follows [47, 17, 21]:
@@ 3 Applicability of DeePC for nonlinear systems
Since the primary focus is on the future behaviors of the lifted state $\Psi$ (which are directly related to the original system states $x$) instead of all the elements of $\Psi_{\chi}$, it is sufficient only to identify matrices $A_{\mathcal{K}}$ and $B_{\mathcal{K}}$. In addition, a projection matrix $D_{\mathcal{K}}$ is used to map the lifted state $\Psi$ back to the original state space. The Koopman-based linear model for nonlinear control system (7) can be described as follows [17, 21, 22, 23]:
@@
where $C_{z}=CD_{\mathcal{K}}$ is the output measurement matrix. (11) characterizes (or more precisely, approximates) the dynamic behaviors of the original nonlinear system in (7) in a higher-dimensional space. It is worth noting that the established Koopman-based LTI model in (11) has the same control input $u$ and output measurements $y$ as the underlying nonlinear system in (7). This implies that for the nonlinear process (7), there exists a higher-dimensional LTI model that can represent or approximate the nonlinear dynamics of the underlying system, with control inputs and measured outputs remaining unaffected by coordinate changes, thereby retaining their original physical meaning.
For any DeePC method that can be applied to the LTI model in (11), it is almost equivalent to apply the same DeePC method to the underlying nonlinear system in (7). On the other hand, based on Willems’ fundamental lemma, if the established Koopman-based LTI model in its lifted space is controllable, a DeePC controller can be formulated following (6). This supports the appropriateness of applying DeePC-based methods that were originally developed for LTI systems, including the conventional DeePC in (6), to the nonlinear system in (7), particularly from a Koopman-based modeling and control perspective.
4 Deep learning-based DeePC control approach
@@
4.1 Structure of the proposed method
A graphical illustration of the proposed Deep DeePC pipeline is presented in Figure 1. Historical open-loop data are utilized to construct a dataset, which is used to train a dense neural network (DNN). This DNN is leveraged to output an approximation of the DeePC operator $g$ in the non-parametric model in (5), based on the constructed training data from the open-loop data. The trajectories of future inputs and system outputs, denoted by $\hat{\mathbf{u}}$ and $\hat{\mathbf{y}}$, will be predicted based on the output of the DNN and (5). The objective is to train the DNN such that future system outputs are driven towards the reference trajectories.
4.2 Deep DeePC
By incorporating the equality constraints $\hat{\mathbf{u}}_{k}=U_{f}g_{k}$ and $\hat{\mathbf{y}}_{k}=Y_{f}g_{k}$ from (6j) into the objective function (6a), the online optimization problem for DeePC in (6) are re-formulated as follows [43]:
@@ 4.2 Deep DeePC
where $\hat{g}_{k}$ is the predicted DeePC operator and $\theta$ includes the trainable parameters of neural network $F_{\theta}$. The inputs to $F_{\theta}$ include the initial input and output trajectories, that is, $\mathbf{u}_{ini,k}$ and $\mathbf{y}_{ini,k}$. Additionally, to enable set-point tracking capability of the proposed Deep DeePC method, the input and output tracking errors between the future and current time instant are also incorporated as inputs to this neural network. Specifically, the input error is the difference between the reference input at current time instant $k$ and the control input at previous time instant $k-1$, i.e., $e_{u,k}=u^{r}_{k}-u_{k-1}$, and the output error is the difference between the reference output at time instant $k+1$ and the latest measured output at time instant $k$, i.e., $e_{y,k}=y^{r}_{k+1}-y_{k}$. The time instants for the input and output errors are different because the system output is one step ahead of the input. The dimension of the inputs to the neural network $F_{\theta}$ is determined based on the selected parameters $T_{ini}$ for DeePC design and the system parameters $n_{u}$ and $n_{y}$, resulting in a dimension of $(n_{u}+n_{y})\times(T_{ini}+1)$. The dimension of the output of $F_{\theta}$ is the same as the dimension of DeePC operator $g_{k}$ in (12), which is $T-T_{ini}-N_{p}+1$.
The objective of this neural network is to output DeePC operator $\hat{g}_{k}$, which can be further used to generate control actions that drive the system towards desired reference outputs – this control objective aligns with that of DeePC (12a). To address the input and output constraints, a soft constraint $\omega(\hat{\mathbf{u}},\hat{\mathbf{y}})$ is incorporated into the objective function. Given a dataset $\mathcal{D}$ composed of multiple $T_{ini}$-step initial input and output trajectories, $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, and their corresponding $N_{p}$-step input and reference outputs, $\mathbf{u}^{r}$ and $\mathbf{y}^{r}$, the objective function used to train the parameters in $\theta$ is as follows:
@@
where $\mathbf{u}_{lb}\in\mathbb{R}^{n_{u}N_{p}}$ and $\mathbf{u}_{ub}\in\mathbb{R}^{n_{u}N_{p}}$ are the lower and upper bounds of the input variables for $N_{p}$ steps, respectively (similarly for $\mathbf{y}_{lb}$ and $\mathbf{y}_{ub}$). $P^{\prime}_{u,lb}=P_{u}\odot M_{u,lb}\in\mathbb{R}^{n_{u}N_{p}\times n_{u}N_{p}}$ and $P^{\prime}_{u,ub}=P_{u}\odot M_{u,ub}\in\mathbb{R}^{n_{u}N_{p}\times n_{u}N_{p}}$ where $P_{u}$ is the weighting matrix of the term that penalizes the violation of input constraint; $M_{u,lb}=\text{diag}([\mathbf{1}_{\hat{\mathbf{u}}(1)<\mathbf{u}_{lb}(1)},\ldots,\mathbf{1}_{\hat{\mathbf{u}}(n_{u}N_{p})<\mathbf{u}_{lb}(n_{u}N_{p})}])$ and $M_{u,ub}=\text{diag}([\mathbf{1}_{\mathbf{u}_{ub}(1)<\hat{\mathbf{u}}(1)},\ldots,$ $\mathbf{1}_{\mathbf{u}_{ub}(n_{u}N_{p})<\hat{\mathbf{u}}(n_{u}N_{p})}])$ are the mask matrices used to penalize the violated terms (similarly for $P^{\prime}_{y,lb}$ and $P^{\prime}_{y,ub}$).
@@
4.3 Training data construction
@@
Training the neural network in (13) based on the objective function in (14) requires a data set that comprises initial input and initial output trajectories (denoted by $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, respectively), and their corresponding future input and reference outputs (denoted by $\mathbf{u}^{r}$ and $\mathbf{y}^{r}$, respectively). The initial input and output trajectories are used as inputs to the neural network in (13), while the future references serve as the labels that are used to compute the values of the objective function, as needed for supervised learning-based training as described in (14). The combined initial trajectories and future references should consist of samples collected at consecutive time instants that span a specific $(T_{ini}+N_{p})$-step time window. In this work, we utilize historical data to construct the input and label data for training the neural network.
Figure 2 provides a graphical illustration of the construction of a training data set based on historical data of the system collected offline. The objective is to construct $(T_{ini}+N_{p})$-step continuous sequences, which can then be divided into input and label data for neural network training. As shown in Figure 2, $(N-T_{ini}-N_{p}+1)$ segments of the $(T_{ini}+N_{p})$-step sequences can be extracted from an $N$-step sequential system trajectory. Within each segment, the first $T_{ini}$ steps of the segment are treated as the initial trajectories $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$ which are the input of the neural network, and the remaining $N_{p}$ steps of the segment are regarded as the future references $\mathbf{u}^{r}$ and $\mathbf{y}^{r}$ corresponding to the initial trajectories which are the labels of the supervised learning. In addition, the input and output errors, $e_{u}$ and $e_{y}$, are computed by the constructed initial trajectories and future references of the selected segment based on (13).
5 Event-based constraint handling
@@
The proposed Deep DeePC in Section 4.2 is highly computationally efficient since it does not require solving online optimization. Meanwhile, it is worth mentioning that the Deep DeePC controller in (17) is not capable of handling hard constraints on either the system inputs or the system outputs. To deal with cases when constraint satisfaction is critical, for example, when safety-related constraints need to be satisfied, we further propose an event-based constraint handling scheme. This scheme can be incorporated into the proposed online optimization-free Deep DeePC (17) to ensure the satisfaction of input and output constraints during online implementation, when it is necessary. An illustrative diagram of the Deep DeePC approach with the constraint handling scheme is presented in Figure 3.
The online implementation of Deep DeePC with the constraint handling scheme is described in Algorithm 1. At each sampling instant $k$, $k\in\mathbb{Z}_{\geq 0}$, the trained DNN model generates the DeePC operator $\hat{g}_{k}$ based on the $\mathbf{u}_{ini,k}$, $\mathbf{y}_{ini,k}$, $e_{u,k}$, $e_{y,k}$ collected during online operation:
@@
    6 Extension of the proposed method
@@
    The online implementation of the well-trained Deep DeePC model for control tasks requires information on the initial input and output trajectories, $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, along with the input and output errors, $e_{u}$ and $e_{y}$. However, in some cases where the steady-state reference inputs corresponding to the set-points are unknown, implementing the current approach becomes impractical. Therefore, an alternative design for Deep DeePC is proposed that does not require steady-state reference inputs.
@@
    where $\tilde{F}_{\tilde{\theta}}$ denotes the neural network without the input of steady-state reference input; $\tilde{\theta}$ denotes the trainable parameters of neural network $\tilde{F}_{\tilde{\theta}}$. The input dimension of the neural network $\tilde{F}_{\tilde{\theta}}$ is reduced to $(n_{u}+n_{y})\times T_{ini}+n_{y}$. Given a data set $\tilde{\mathcal{D}}$ that composed of multiple $\mathbf{u}_{ini}$, $\mathbf{y}_{ini}$, and corresponding $\mathbf{y}^{r}$, the objective function used to train the neural network $\tilde{F}_{\tilde{\theta}}$ is:
@@
    7 Case study on gene regulatory network
    7.1 Process description

### v3

@@ Introduction
The Koopman theory-based framework [16] holds great promise for data-driven modeling and control of nonlinear systems. By constructing a linear model within a lifted state-space for a general nonlinear system from offline data, Koopman modeling facilitates the application of linear control theory to nonlinear systems and processes. Notable advancements include Koopman-based model predictive control (MPC) approaches [17, 18, 20, 19, 21], which develop linear MPC schemes for nonlinear systems, thereby maintaining convex online optimization despite the inherent nonlinearity of the underlying systems and processes. Koopman-based modeling and control have been widely applied across various fields, including industrial processes [22, 23, 24, 25], vehicles and robotics [26, 27, 28], and power systems [29, 30, 31]. The development of Koopman MPC typically requires full-state feedback from the systems, with a few exceptions [32, 33]. An input-output Koopman control method was proposed, where the Koopman-based controller only requires measurements of partial state variables directly used for calculating operational costs [32]. A non-exact multi-step Koopman predictor was identified using only input and output data for the predictive control of nonlinear systems [33]. We note that, in practical applications, measuring certain state variables can be challenging or costly. The need for full-state measurements poses significant challenges to the broader applications of Koopman-based control methods in complex industrial systems.
@@
In recent years, data-enabled predictive control (DeePC) has received increasing attention for its ability to address constrained optimal control using only input and output data [36, 34, 35]. DeePC leverages Willems’ fundamental lemma [37] to form a non-parametric representation of the system using a Hankel matrix established based on pre-collected input and output trajectories. Willems’ fundamental lemma indicates that any input and output trajectories of a linear system are spanned by the Hankel matrix when the pre-collected input trajectory is persistently exciting [37]. DeePC produces the optimal control action by solving an optimization problem in a receding-horizon manner, which is similar to MPC [38]. In the optimization problem, the number of the decision variables is dependent on the dimension of the Hankel matrix. To satisfy the persistently exciting condition, a sufficiently large amount of historical data is typically needed. This can lead to relatively large dimensions for both the Hankel matrix and the decision variables, which will increase the computational complexity of the online optimization problem associated with DeePC.
@@
In the existing literature, efforts have been focused on addressing the computational challenges of DeePC by reducing the size or complexity of the Hankel matrix and optimization problems. Singular value decomposition was leveraged on collected data to reduce the dimensionality of the DeePC problem, thus decreasing computational complexity [39]. Proper orthogonal decomposition was implemented to formulate a reduced order control problem, which was subsequently solved using a customized MPC solver [40]. LQ factorization was employed to redefine the decision variables, replacing the Hankel matrix with a lower triangular matrix [41, 42]. Lagrange multiplier was used to obtain a nominal solution to the DeePC optimization problem, and an adaption approach was proposed to efficiently update the nominal solution without recomputing the optimal solution [43]. The primal-dual algorithm was leveraged to solve the DeePC problem recursively, and the fast Fourier transform was used to expedite the computation of Hankel matrix-vector products [44]. A size-invariant differentiable convex problem was proposed to learn the scoring function of the DeePC problem recursively, where the scoring function evaluates the likelihood of the predicted input and output trajectories belonging to the system [45]. While these existing methods improve the computational feasibility of DeePC, they still involve solving optimization problems or iterating algorithms during real-time applications, highlighting ongoing challenges in achieving real-time efficiency.
@@
    We propose a deep learning-enabled DeePC method, referred to as Deep DeePC, to improve online computational efficiency; this method does not need to solve any online optimization when constraints are either absent or not violated.
@@
    We use two benchmark examples to evaluate the performance of the proposed approach. Good control performance is achieved, system constraints are satisfied, and the computation time is significantly reduced as compared to conventional DeePC.
Preliminaries and problem formulation
Notation
$\mathbb{R}$ denotes the set of real numbers. $\mathbb{Z}_{>0}$ and $\mathbb{Z}_{\geq 0}$ denote the sets of positive integers and the set of non-negative integers, respectively. $\mathbb{E}$ denotes the expectation. $\|x\|_{Q}^{2}$ is the square of the weighted Euclidean norm of vector $x$ with positive-definite weighting matrix $Q$, computed as $\|x\|_{Q}^{2}:=x^{\top}Qx$. $\text{diag}(\cdot)$ denotes a diagonal matrix. $I_{n}$ is an identity matrix of dimension $n$. $\odot$ denotes the Hadamard product. $\mathbf{1}_{a<b}$ denotes an indicator function; it equals $1$ if $a<b$, and equals 00 otherwise. $\mathcal{U}(a,b)$ denotes a uniform distribution with lower and upper bounds $a$ and $b$, respectively. $x(i)$ denotes the $i$th variable of state vector $x$. $x_{k}\in\mathbb{R}^{n_{x}}$ is the state vector at time instant $k$, and $\{x\}_{j}^{l}:=[x_{j}^{\top},\ldots,x_{l}^{\top}]^{\top}$ contains the state sequence from time instant $j$ to $l$.
Non-parametric representation of linear systems
@@
Let $L$, $T\in\mathbb{Z}_{>0}$ and $T\geq L$. Let $\mathbf{u}_{T}^{d}:=\{u^{d}\}_{1}^{T}\in\mathbb{R}^{n_{u}T}$ and $\mathbf{y}_{T}^{d}:=\{y^{d}\}_{1}^{T}\in\mathbb{R}^{n_{y}T}$ denote the input and output sequences for $T$ time instants, respectively. The superscript $d$ indicates the corresponding data are historical data. For the input sequence $\mathbf{u}_{T}^{d}$, the Hankel matrix of depth $L$ is defined as follows:
@@
Data-enabled predictive control (DeePC)
@@
$x_{k+1}$ $=f(x_{k},u_{k})$ (7)
$y_{k}$ $=Cx_{k}$
where $x_{k}\in\mathbb{X}\subset\mathbb{R}^{n_{x}}$ is the system state vector; $u_{k}\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$ is the control input vector; $y_{k}\in\mathbb{Y}\subset\mathbb{R}^{n_{y}}$ is the output vector; $f:\mathbb{X}\times\mathbb{U}\to\mathbb{X}$ is a nonlinear function that characterizes the state dynamics of the process; $C\in\mathbb{R}^{n_{y}\times n_{x}}$ is the system matrix; $\mathbb{X}$, $\mathbb{U}$, and $\mathbb{Y}$ are compact sets. We consider nonlinear systems in which the controlled outputs linearly depend on the state vectors. Our objective is to address the control for the nonlinear processes using only input and output data in an efficient manner within the DeePC framework.
@@
Applicability of DeePC for nonlinear systems
We will use the conventional DeePC described in (6) as the foundation of developing our deep learning-enabled DeePC method. Given that (6) was originally developed for LTI systems based on Lemma 1, in this section, we leverage the concept of Koopman modeling for control to justify the suitability of using conventional DeePC in (6) as the basis for developing DeePC-based controllers for nonlinear systems in (7).
Consider a discrete-time nonlinear control system described in (7). Based on the Koopman theory for controlled systems [17], the state vector can be extended to include control inputs, that is, $\chi_{k}=[x_{k}^{\top},u_{k}^{\top}]^{\top}$. Accordingly, there exists an infinite-dimensional nonlinear lifting mapping $\Psi_{\chi}$ that allows the dynamics of the lifted state to be governed by a Koopman operator $\mathcal{K}$ as follows [47, 17, 21]:
@@
Since the primary focus is on the future behaviors of the lifted state $\Psi$ (which are directly related to the original system state $x$) instead of all the elements of $\Psi_{\chi}$, it is sufficient only to identify matrices $A_{\mathcal{K}}$ and $B_{\mathcal{K}}$. In addition, a projection matrix $D_{\mathcal{K}}$ is used to map the lifted state $\Psi$ back to the original state space. The Koopman-based linear model for nonlinear control system (7) can be described as follows [17, 21, 22, 23]:
@@
where $C_{z}=CD_{\mathcal{K}}$ is the output matrix. (11) characterizes (or more precisely, approximates) the dynamic behaviors of the original nonlinear system in (7) in a higher-dimensional space.
It is worth noting that the established Koopman-based LTI model in (11) has the same control input $u$ and controlled output $y$ as the underlying nonlinear system in (7). This implies that for the nonlinear system in (7), there exists a higher-dimensional LTI model that can represent or approximate the nonlinear dynamics of the underlying system, with control inputs and measured outputs remaining unaffected by coordinate changes, thereby retaining their original physical meaning. Meanwhile, the conventional DeePC method, originally developed for LTI systems in (6), may be directly applied to systems described by the Koopman-based model in (11) based on control input $u$ and controlled output $y$. Given that it is feasible to construct an accurate Koopman model in (11) as a high-fidelity surrogate for the nonlinear system in (7), the DeePC scheme developed based on (11) can serve as an effective predictive control solution for the nonlinear system in (7). The above analysis supports the suitability of applying DeePC for control of the nonlinear system in (7).
Deep learning-based DeePC control approach
@@
Structure of the proposed method
A graphical illustration of the proposed Deep DeePC pipeline is presented in Figure 1. Historical open-loop data are utilized to construct a dataset, which is used to train a dense neural network (DNN). This DNN is used to output an approximation of the DeePC operator $g$ in (6). The trajectories of future input and system output, denoted by $\hat{\mathbf{u}}$ and $\hat{\mathbf{y}}$, will be predicted based on the output of the DNN and (5). The objective is to train the DNN such that future system outputs are driven towards the reference trajectories.
Deep DeePC
By incorporating the equality constraints $\hat{\mathbf{u}}_{k}=U_{f}g_{k}$ and $\hat{\mathbf{y}}_{k}=Y_{f}g_{k}$ from (6j) into the objective function (6a), the online optimization problem for DeePC in (6) is re-formulated as follows [43]:
@@
where $\hat{g}_{k}$ is the predicted DeePC operator, $\theta$ includes the trainable parameters of neural network $F_{\theta}$, $e_{u,k}=u^{r}_{k}-u_{k-1}$ is the input tracking error that represents the difference between the reference input at current instant $k$ and the most recent control input at time instant $k-1$, and $e_{y,k}=y^{r}_{k+1}-y_{k}$ is the output tracking error that represents the difference between the reference output at time instant $k+1$ and the latest measured output at time instant $k$. The time instants for the input and output tracking errors are different because the system output is one step ahead of the input. The inputs to $F_{\theta}$ include the initial input and output trajectories, that is, $\mathbf{u}_{ini,k}$ and $\mathbf{y}_{ini,k}$, which contain information for the past $T_{ini}$ steps. Since the neural network is used to generate future control inputs for set-point tracking, $e_{u,k}$ and $e_{y,k}$, which contain information for the most recent input and output values and their corresponding future references for the next step, are also incorporated as inputs to this neural network. The dimension of the inputs to the neural network $F_{\theta}$ is determined based on the selected parameters $T_{ini}$ for DeePC design and the system parameters $n_{u}$ and $n_{y}$, resulting in a dimension of $(n_{u}+n_{y})\times(T_{ini}+1)$. The dimension of the output of $F_{\theta}$ is the same as the dimension of DeePC operator $g_{k}$ in (12), which is $T-T_{ini}-N_{p}+1$.
The objective of this neural network is to output DeePC operator $\hat{g}_{k}$, which can be further used to generate control actions that drive the system towards desired reference outputs – this control objective aligns with that of DeePC (12a). To address the input and output constraints, a soft constraint $\omega(\hat{\mathbf{u}},\hat{\mathbf{y}})$ is incorporated into the objective function. Given a dataset $\mathcal{D}$ composed of multiple $T_{ini}$-step initial input and output trajectories, $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, and their corresponding $N_{p}$-step reference input and reference output trajectories, $\mathbf{u}^{r}$ and $\mathbf{y}^{r}$, the objective function used to train the parameters in $\theta$ is as follows:
@@
where $\mathbf{u}_{lb}\in\mathbb{R}^{n_{u}N_{p}}$ and $\mathbf{u}_{ub}\in\mathbb{R}^{n_{u}N_{p}}$ are the lower and upper bounds of the input variables for $N_{p}$ steps, respectively (similarly for $\mathbf{y}_{lb}$ and $\mathbf{y}_{ub}$). $P^{\prime}_{u,lb}=P_{u}\odot M_{u,lb}\in\mathbb{R}^{n_{u}N_{p}\times n_{u}N_{p}}$ and $P^{\prime}_{u,ub}=P_{u}\odot M_{u,ub}\in\mathbb{R}^{n_{u}N_{p}\times n_{u}N_{p}}$, where $P_{u}$ is the weighting matrix of the term that penalizes the violation of input constraint; $M_{u,lb}=\text{diag}([\mathbf{1}_{\hat{\mathbf{u}}(1)<\mathbf{u}_{lb}(1)},\ldots,\mathbf{1}_{\hat{\mathbf{u}}(n_{u}N_{p})<\mathbf{u}_{lb}(n_{u}N_{p})}])$ and $M_{u,ub}=\text{diag}([\mathbf{1}_{\mathbf{u}_{ub}(1)<\hat{\mathbf{u}}(1)},\ldots,$ $\mathbf{1}_{\mathbf{u}_{ub}(n_{u}N_{p})<\hat{\mathbf{u}}(n_{u}N_{p})}])$ are the mask matrices used to penalize the violated terms (similarly for $P^{\prime}_{y,lb}$ and $P^{\prime}_{y,ub}$).
@@
Remark 1
In the existing literature, neural networks have been utilized to approximate the control policy of MPC controllers [49, 50], which is another effective framework to substantially increase the online computation speed of optimal control. This type of approach typically relies on an existing MPC controller developed based on a dynamic model to generate closed-loop data for neural network training. However, when a dynamic model is unavailable and only input and output data are accessible, the proposed Deep DeePC method can serve as a viable alternative.
Training data construction
@@
Training the neural network in (13) based on the objective function in (14) requires a dataset that comprises initial input and initial output trajectories (denoted by $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, respectively), and their corresponding future reference input and reference output trajectories (denoted by $\mathbf{u}^{r}$ and $\mathbf{y}^{r}$, respectively). The initial input and output trajectories are used as inputs to the neural network in (13), while the future references serve as the labels that are used to compute the values of the objective function, as needed for supervised learning-based training as described in (14). The combined initial trajectories and future references should consist of samples collected at consecutive time instants that span a specific $(T_{ini}+N_{p})$-step time window. In this work, we utilize historical data to construct the input and label data for training the neural network.
Figure 2 provides a graphical illustration of the construction of a training dataset based on historical data of the system collected offline. The objective is to construct $(T_{ini}+N_{p})$-step continuous sequences, which can then be divided into input and label data for neural network training. As shown in Figure 2, $(N-T_{ini}-N_{p}+1)$ segments of the $(T_{ini}+N_{p})$-step sequences can be extracted from an $N$-step sequential system trajectory. Within each segment, the first $T_{ini}$ steps of the segment are treated as the initial trajectories $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, which are the inputs of the neural network, and the remaining $N_{p}$ steps of the segment are regarded as the future references $\mathbf{u}^{r}$ and $\mathbf{y}^{r}$ corresponding to the initial trajectories, which are the labels of the supervised learning. In addition, the input and output errors, $e_{u}$ and $e_{y}$, are computed by the constructed initial trajectories and future references of the selected segment based on (13).
Event-based constraint handling
@@
The proposed Deep DeePC approach in the “Deep DeePC” section is highly computationally efficient since it does not require solving online optimization. Meanwhile, it is worth mentioning that the Deep DeePC controller in (17) is not capable of handling hard constraints on either the system inputs or the system outputs. To deal with cases when constraint satisfaction is critical, for example, when safety-related constraints need to be satisfied, we further propose an event-based constraint handling scheme. This scheme can be incorporated into the proposed online optimization-free Deep DeePC (17) to ensure the satisfaction of input and output constraints during online implementation, when it is necessary. An illustrative diagram of the Deep DeePC approach with the constraint handling scheme is presented in Figure 3.
The online implementation of Deep DeePC with the constraint handling scheme is described in Algorithm 1. At each sampling instant $k$, $k\in\mathbb{Z}_{\geq 0}$, the trained DNN model generates the DeePC operator $\hat{g}_{k}$ based on the $\mathbf{u}_{ini,k}$, $\mathbf{y}_{ini,k}$, $e_{u,k}$, and $e_{y,k}$ collected during online operation:
@@
    Extension of the proposed method
@@
    The online implementation of the well-trained Deep DeePC model for control tasks requires information on the initial input and output trajectories, $\mathbf{u}_{ini}$ and $\mathbf{y}_{ini}$, along with the input and output errors, $e_{u}$ and $e_{y}$. However, in some cases where the steady-state reference inputs corresponding to the set-points are unknown, implementing the current approach becomes impractical. Therefore, an alternative design for Deep DeePC that does not require steady-state reference inputs is proposed.
@@
    where $\tilde{F}_{\tilde{\theta}}$ denotes the neural network without requiring the steady-state reference input; $\tilde{\theta}$ denotes the trainable parameters of neural network $\tilde{F}_{\tilde{\theta}}$. The input dimension of the neural network $\tilde{F}_{\tilde{\theta}}$ is reduced to $(n_{u}+n_{y})\times T_{ini}+n_{y}$. Given a dataset $\tilde{\mathcal{D}}$ that is composed of multiple $\mathbf{u}_{ini}$, $\mathbf{y}_{ini}$, and corresponding $\mathbf{y}^{r}$, the objective function used to train the neural network $\tilde{F}_{\tilde{\theta}}$ is defined as follows:
@@
    Case study on gene regulatory network
    Process description